# Impact chronology of leftover planetesimals

### v1

@@ 4 Spherule layer constraints on the Archean impact rate and source population
What about the E-belt of Bottke et al. (2012) being the source? Brasser et al. (2020) conducted dynamical simulations of the E-belt for 1 Gyr. For this work these were extended to 2 Gyr. For the E-belt the impact probability with Earth is , requiring 340k bodies to leak from the E-belt between 3.47 Ga and 2 Ga to supply all of the terrestrial impacts. The fraction remaining can be fit by , so that at late times Myr-1. In the simulations the E-belt loses an additional 93% of its mass between 3.47 Ga and 2 Ga, so that I can safely assume that at 3.47 Ga the E-belt had 340k objects with km. The simulation output shows that at 3.47 Ga the E-belt only has 3.9% of its mass at 4.5 Ga, implying that the primordial number of E-belt objects with km is approximately 8.7M assuming no collisional evolution. The required mass for this many objects is about 0.002 , which is a factor of a few higher than what was suggested by Bottke et al. (2012). It is therefore more likely that the Archean impacts that created the terrestrial spherule beds were predominantly () supplied by leftover planetesimals on meta-stable orbits slowly leaking onto Earth-crossing orbits from beyond Mars.
@@ 5 Conclusions
I have simulated the dynamical evolution of leftover planetesimals for 1 billion years from three different models of terrestrial planet formation: the Grand Tack (Walsh et al., 2011), the Depleted Disc (Izidoro et al., 2014), and implantation (Brasser, 2025). Each one of these models yields a different initial distribution of orbital elements for the leftover planetesimals. In particular, the initial fraction of Earth crossers for the GT is twice as high as for DD, which has implications for the dynamical evolution and impact rate onto the Earth and Mars. The DD model leaves too much mass after 1 Gyr to be consistent with the current mass of the main asteroid belt.
I fit the cumulative number of impacts and the remaining fraction of planetesimals with a function that is a sum of three to four exponentials whose e-folding timescales cluster around 10 Myr, 35 Myr, 100 Myr and 200 Myr. These distinct timescales are the result of different processes removing the planetesimals at different rates, with high-eccentricity Earth crossers and planetesimals in the secular resonance being removed first (), followed by Earth crossers (), Mars crossers () and finally objects leaking on to Mars crossers from the E-belt and the inner main belt (). The timescale is of the same order as that of the lunar chronology of Neukum et al. (2001) and Werner et al. (2023), suggesting that the lunar craters are sourced predominantly from an ancient population of Mars crossers, and that the number of Earth crossers at the time that the Moon formed had to be rather low. As such, the initial perihelion distribution of the leftover planetesimals plays a decisive role in the overall rate of decline of the population and the impact rates onto the Earth and Mars. These results place constraints on models of terrestrial planet formation, wherein the number of Earth crossers at the time of the Moon’s formation had to be relatively low compared to the Mars crossers.
To calibrate the impact rate at late times I use the Archean spherule beds as listed in Johnson & Melosh (2012), converting the average diameter of the impactor to the number of objects with diameter km striking the Earth over 1.47 Gyr between 3.47 Ga and 2 Ga. I conclude that the most likely source of these impacts are the leftover planetesimals rather than the main belt or the E-belt, and that this Archean flux requires a primordial leftover population with a mass of about 0.015 at 4.5 Ga. Implications for the lunar cratering rate and the amount of mass accreted by the Moon and Mars will be presented in the future.
@@ Acknowledgements.
I thank Stephanie Werner and Emily W. Wong for comments. I acknowledge EuroHPC Joint Undertaking for awarding the project EHPC-REG-2024R01-093 access to Vega at IZUM, Slovenia. This study is supported by the Research Council of Norway through its Centres of Excellence funding scheme, project No. 332523 PHAB. GENGA can be obtained from https://bitbucket.org/sigrimm/genga/.References

### v3

@@ 4 Spherule layer constraints on the Archean impact rate and source population
What about the E-belt of Bottke et al. (2012) being the source? Brasser et al. (2020) conducted dynamical simulations of the E-belt for 1 Gyr. For this work these were extended to 2 Gyr. For the E-belt the impact probability with Earth is , requiring 340k bodies to leak from the E-belt between 3.47 Ga and 2 Ga to supply all of the terrestrial impacts. The fraction remaining can be fitted as , so that at late times Myr-1. In the simulations the E-belt loses an additional 93% of its mass between 3.47 Ga and 2 Ga, so that I can safely assume that at 3.47 Ga the E-belt had 340k objects with km. The simulation output shows that at 3.47 Ga the E-belt only has 3.9% of its mass at 4.5 Ga, implying that the primordial number of E-belt objects with km is approximately 8.7M assuming no collisional evolution. The required mass for this many objects is about 0.002 , which is a factor of a few higher than what was suggested by Bottke et al. (2012). It is therefore more likely that the Archean impacts that created the terrestrial spherule beds were predominantly () supplied by leftover planetesimals on meta-stable orbits slowly leaking onto Earth-crossing orbits from beyond Mars.
@@ 5 Conclusions
I have simulated the dynamical evolution of leftover planetesimals for 1 billion years from three different models of terrestrial planet formation: the GT (Walsh et al., 2011), the DD (Izidoro et al., 2014), and ABI (Brasser, 2025). Each one of these models yields a different initial distribution of orbital elements for the leftover planetesimals. In particular, the initial fraction of Earth-crossers for the GT is twice as high as for DD, which has implications for the dynamical evolution and impact rate onto the Earth and Mars. The DD model leaves too much mass after 1 Gyr to be consistent with the current mass of the main asteroid belt.
I fitted the cumulative number of impacts and the remaining fraction of planetesimals with a function that is a sum of three to four exponentials whose e-folding timescales cluster around 10 Myr, 35 Myr, 100 Myr, and 200 Myr. These distinct timescales are the result of different processes removing the planetesimals at different rates, with high-eccentricity Earth-crossers and planetesimals in the secular resonance being removed first (), followed by Earth-crossers (), Mars-crossers (), and finally objects leaking onto Mars-crossing orbits from the E-belt and the inner main asteroid belt (). The timescale is of the same order as that of the lunar chronology of Neukum et al. (2001) and Werner et al. (2023), suggesting that the lunar craters are sourced predominantly from an ancient population of Mars-crossers, and that the number of Earth-crossers at the time that the Moon formed had to be rather low. As such, the initial perihelion distribution of the leftover planetesimals plays a decisive role in the overall rate of decline of the population and the impact rates onto the Earth and Mars. These results place constraints on models of terrestrial planet formation, wherein the number of Earth-crossers at the time of the Moon’s formation had to be lower than Mars-crossers.
To calibrate the impact rate at late times, I used the Archean spherule beds as listed in Johnson & Melosh (2012), converting the average diameter of the impactor to the number of objects with diameter km striking the Earth over 1.47 Gyr between 3.47 Ga and 2 Ga. I conclude that the most likely source of these impacts are the leftover planetesimals rather than the main belt or E-belt asteroids, and that this Archean flux requires a primordial leftover population with a mass of about 0.015 at 4.5 Ga. Implications for the lunar cratering rate and the amount of mass accreted by the Moon and Mars will be presented in a future paper.
@@ Acknowledgements.
I thank Stephanie Werner and Emily W. Wong for comments, and John Chambers for his constructive review. I acknowledge EuroHPC Joint Undertaking for awarding the project EHPC-REG-2024R01-093 access to Vega at IZUM, Slovenia. This study is supported by the Research Council of Norway through its Centres of Excellence funding scheme, project No. 332523 PHAB. GENGA can be obtained from https://bitbucket.org/sigrimm/genga/.References
@@
- Neukum et al. (1975) Neukum, G., Koenig, B., & Arkani-Hamed, J. 1975, Moon, 12, 201
@@
- Valsecchi et al. (1997) Valsecchi, G. B., Froeschlé, C., & Gonczi, R. 1997, Planet. Space Sci., 45, 1561
- Valsecchi et al. (2003) Valsecchi, G. B., Milani, A., Gronchi, G. F., & Chesley, S. R. 2003, A&A, 408, 1179